OpenStudy (anonymous):

antiderivative(18)/(5-x) = kln[5-x]+c Find the exact value of k that makes the antidifferentiation formula true.

OpenStudy (anonymous):

i got 18 but is wrong

OpenStudy (anonymous):

@ jorourk3 \[\int\limits_{}^{}\frac{ 18 }{ 5-x }=-18\ln(x-5)+C\]=kln[5-x]+c seems like k=-1 ? yes or no?

OpenStudy (anonymous):

im sorry its -18 :D

OpenStudy (anonymous):

k would be -18 then by what your saying...wouldnt it? that is also absolute value signs around ln[x-5] btw...does that change it from what you are saying?

OpenStudy (anonymous):

its correct! thanks! i dont understand why it is negative though?

OpenStudy (anonymous):

its negative bec you let u=5-x then du= -dx

OpenStudy (anonymous):

or du/dx=-1

OpenStudy (anonymous):

ok thanks....can you help me with a similar one please? :)

OpenStudy (anonymous):

ok yw ..:D just post them again so that others can see them